OpenStudy (anonymous):

Given that f(x) = 4x - 3 and g(x) =solve for g(f(2))

OpenStudy (freckles):

Well plug 2 in to the f function

OpenStudy (freckles):

then take that result in plug it into the g function

OpenStudy (freckles):

so first f(2)=4(2)-3=?

OpenStudy (anonymous):

5

OpenStudy (freckles):

\[g(f(2))=g(4 \cdot 2-3)=g(8-3)=g(5)\]

OpenStudy (freckles):

now find g(5) by using whatever g(x) equals

OpenStudy (anonymous):

Oh wait the equation is wrong hold on!

OpenStudy (anonymous):

Given that f(x) = 4x - 3 and g(x) = 2x-1/3 solve for g(f(2)).

OpenStudy (freckles):

g(x)=(2x-1)/3 find g(5) by replacing the x's with 5's

OpenStudy (anonymous):

3?

OpenStudy (freckles):

yes [2(5)-1]/3=[10-1]/3=9/3=3

OpenStudy (anonymous):

And thats the answer?

OpenStudy (anonymous):

@freckles